OpenStudy (anonymous):

When completing the square on the equation c² + 11c = 12, the resulting equation is:

OpenStudy (jim766):

take 11 and divide by 2

OpenStudy (anonymous):

and then what else ?

OpenStudy (jim766):

then square it and add it to both sides of the equation

OpenStudy (anonymous):

okay an what would, \[6m ^{2}+5-3=0\] look like when i square it...there is a \[\pm \] in it

OpenStudy (jim766):

is it 6m^2 + 5m - 3 = 0 ?

OpenStudy (anonymous):

yes sir !

OpenStudy (jim766):

6m^2 + 5m - 3 = 0 +3 + 3 6m^2 + 5m = 3 then divide everything by 6

OpenStudy (jim766):

m^2 + 5/6x = 1/2 then take 5/6 div by 2 and square and add that to both sides

OpenStudy (jim766):

I would never do this one this way...if you have a choice..